Please help me solve this equation:4x+32y=24
Answer by cynthia.perez(4)

you would have to see which one you want to solve for.
lets say your solving for x.
4x+32y=24
subtract 32y
4x=24-32y
divide by 4
x=6-8y
you the substitute for x.
4(6-8y)+32y=24
24-32y+32y=24
you would have to combine the Y's because thy are on the same side. you get 24=24. that i correct.
hoped this helped.